Let f(x)= -5x-1 and g(x)= x^2+4
Find (fog)(-1)

Answer by Aldorozos(172)

(fog)(-1) = -5(x^2 + 4) -1 here all we have done is we have replace x in f(x) with x^2 +4
(fog)(-1) = -5(x^2 + 4) -1 now we have to replace x with -1
= -5[(-1)^2 +4] -1= -5(1+4)-1